# Public Infrastructure Investments for Space Market Development

### v1

@@ 3 An application to commercial crewed space stations
To estimate the costs of these two architectures, I derive cost data from comparable space station components currently under development or currently operational. Cost estimates for core module elements come from NASA’s Gateway program, which provides the closest available analogs for long-duration crewed space infrastructure: the Power and Propulsion Element (PPE) contract provides cost data for power and thermal systems, the Habitation and Logistics Outpost (HALO) contract provides data for pressurized volume and life support integration, and the Canadian Space Agency’s (CSA) Canadarm3 contract provides data for robotic capabilities. Habitat module construction costs derive from the European Space Agency’s (ESA) Columbus laboratory module, while operating costs are estimated using Columbus’ share of total ISS habitable volume applied to current ISS operating expenses. All cost figures are adjusted to 2025 dollars using NASA’s New Start Inflation Index, which accounts for space technology-specific cost escalation. [^12] I exclude several cost categories from this analysis. Insurance costs, which can be significant for crewed systems, are omitted due to indications that private insurance markets lack capacity to cover fully privately owned crewed space stations (MacDonald et al., 2024). Launch costs for initial system deployment are excluded, though including them would favor the shared core architecture. Training costs and supply-side structure details for crew operations are abstracted away, though transportation monopolies or inadequate training facilities could complicate the economics of human spaceflight. Integration costs for the core module are assumed to be 15% of total element costs—an assumption I make for analytical transparency. The resulting cost estimates provide a foundation for comparing the economic implications of different infrastructure sharing approaches.
Figure 2 shows an example of a core module, Figure 3 shows an example of a robotic arm, and Figure 4 shows an example of a habitat module.
@@
Table 2 shows a list of core elements comprising a core module as described and, by analogy to roughly similar systems, approximate costs. Table 3 shows the costs for a habitat module. All figures are expressed in millions of 2025 U.S. dollars and rounded to the nearest million.
Table 2: Core module element cost breakdown and annual costs
@@
Table 3: Habitat module cost breakdown and annual costs
@@
Operations cost \$267 M/year Columbus has total volume of 75 m3 (ESA, 2025b). ISS has total habitable volume of 388 m3 (NASA, 2025b). 2021 NASA OIG found ISS operating expenses steady at roughly \$1.2B/year over 2016-2020 (NASA, 2021b). FY2026 NASA President’s Budget Request projects similar costs from 2025-2030 (NASA, 2025a). Estimated as ISS total operating expenses scaled by Columbus’ share of habitable volume: $(75/388)\times 1200$.
Habitat module total cost annuity \$417 M/year
These cost estimates allow direct application of the Section 2 framework to assess how the choice between independent stations and shared infrastructure affects competitive viability. Both architectures face identical expected economic conditions: \$500M annually in market demand split between competitors, and \$1B in total capturable program budget. The architectures differ in how this budget gets allocated between direct purchases from firms versus investment in shared infrastructure. In the independent free-flyer architecture, the program budget flows entirely to direct purchases (\$500M per firm assuming two competitors). In the shared core architecture, the program invests in infrastructure—life support, power, docking, and robotic capabilities—that costs \$509M annually based on the component estimates in Table 2, leaving \$491M for direct purchases of habitat services (\$246M per firm). I assume firms value the shared core functions identically whether provided by public infrastructure or developed independently. [^13] This reflects the functional equivalence of core capabilities across architectures. The shared core provides the same life support, power, and docking functions that firms would otherwise develop independently. The \$509M annual value represents what firms would need to spend to develop equivalent capabilities, not necessarily the agency’s actual financing costs, which might differ due to existing assets, different financing terms, or economies of scale in procurement. Under this assumption, firms in the shared core architecture benefit from \$509M in cost reduction while receiving \$246M in direct agency purchases. Table 4 presents the resulting economic parameters for both architectures, while Figure 5 maps these onto the sustainable competition diagram to show how well each approach can sustain competition.
Table 4: Economic parameters under alternative space station architectures
@@
Market revenue ($R^{M}_{i}$) \$250 M/year/firm \$250 M/year/firm
@@
Table 4 shows that under the assumed conditions, the independent free-flyer architecture generates negative industry profits of \$355M annually; two competing firms cannot both earn positive returns. Individual firms in this scenario face \$927M in annual costs while receiving only \$750M in combined government and market revenues (\$500M from government purchases plus \$250M from the market), resulting in \$177M annual losses per firm. The shared core architecture, by contrast, generates positive industry profits of \$154M annually, with each firm earning \$77M in annual profits after accounting for the \$509M cost reduction from shared infrastructure. Figure 5 maps these results onto the sustainable competition diagram. The free-flyer architecture falls into the region where only monopoly provision is economically viable, while the shared core architecture can sustain two firms. Under the shared core approach, firms earn profits available for reinvestment or dividends, though not enough to incent entry of additional competitors.
@@ 3.1 Exploring alternative assumptions
The baseline analysis demonstrates how shared infrastructure can enable competition under a given set of conditions. The framework can also be used to explore how different assumptions affect economic architecture. Two economic variables deserve particular attention because they lie largely outside space agency planners’ control yet can significantly influence market development prospects: the scale of market demand for space station services, and the return expectations of private investors. Market demand reflects the broader commercial ecosystem for space-based research, manufacturing, and other applications—markets that may grow rapidly if early demonstrations prove successful, or remain limited if technical or economic barriers persist. Investor return expectations, meanwhile, fluctuate with broader macroeconomic conditions including interest rates, risk appetites, and alternative investment opportunities. By examining how the sustainable competition diagram changes under alternative values for these parameters, planners can assess the robustness of different economic architectures and identify conditions under which architectures might succeed.
@@
Figure 6 shows how doubling market demand from \$500M to \$1B annually affects both architectures. This scenario reflects an optimistic case where early space station demonstrations catalyze broader commercial adoption across research institutions, pharmaceutical companies, manufacturing firms, and other potential users. Under these conditions, both architectures become more viable. The independent free-flyer architecture, which generated \$355M in annual losses under baseline conditions, now produces positive industry profits and can sustain two competitors with margin for reinvestment or dividends. Each firm in the free-flyer scenario would earn \$500M annually from non-government customers plus \$500M from government purchases, covering the \$927M total costs with \$73M profit margin per firm. The shared core architecture performs even better, generating sufficient profits to sustain four competing firms rather than the two supported under baseline demand, with enough profits to provide returns to existing investors and potentially incent entry of additional competitors. Note, however, new entrants would not be able to enter on the same terms as incumbents without more capacity on the shared core or a new core. Further analysis of this scenario can identify threshold levels of market demand such that direct purchases alone may sustain competitive markets without requiring shared infrastructure. However, this outcome depends on successfully developing the commercial ecosystem—an uncertain prospect that may take years or decades to materialize, if ever.
@@
Figure 6: Sustainable competition diagram showing the relationship between industry-level direct government purchases, industry-level total costs, and the potential number of profitable firms for the crewed space stations example with market demand of \$1000 M/year. Shaded regions indicate the expected sustainable number of competitors under different system cost and government-derived revenues.
@@
Figure 7 shows how doubling investor return expectations from 5% to 10% annually affects both architectures. This scenario reflects tighter macroeconomic conditions where rising interest rates, increased risk premiums, or competition from alternative investments raise the hurdle rate for space technology projects. Under 10% return requirements, neither architecture can sustain two competing firms. The independent free-flyer architecture, already unprofitable for two firms under baseline conditions, becomes even less viable with higher costs of capital. The shared core architecture, which supported two firms under baseline conditions, also falls into the monopoly-only region. If there were two firms in the shared core architecture, both would require higher annual profits to compensate investors for the increased return requirement. But revenues and infrastructure-driven cost reductions remain unchanged, and expected cashflows that allowed firms to raise capital under baseline conditions no longer appear as attractive.
@@
These scenarios reveal several practical insights for public space agencies seeking to develop space markets. First, the choice between shared infrastructure and direct purchases can depend on economic conditions that are outside the agency’s control. Under favorable conditions—e.g., strong market demand and low capital costs—direct purchases may suffice to sustain competition without requiring shared infrastructure investments. Under challenging conditions—e.g., limited market demand or tight capital markets—even well-designed shared infrastructure may prove insufficient to maintain competitive supply. Second, given how external economic factors can dominate program design choices in determining market outcomes, timing market development efforts to coincide with favorable economic conditions may be as important as program design choices. Third, the framework provides a systematic way to assess these trade-offs and communicate them to stakeholders, but depends on credible estimates of costs, market demand, and competitive dynamics that remain challenging to develop for novel space technology systems. The estimates may evolve with technical and economic conditions, requiring ongoing assessment and planning.
4 Future Research Directions
I have demonstrated the use of a simple framework to map the number of firms that can profitably provide goods and services derived from a space technology system under different public funding and industry cost structures. The framework demonstrates the multiplier effect of non-rival benefits from shared infrastructure and the sensitivity of market development programs to non-technological factors like market demand expectations and investor return requirements. Though highly stylized, the framework’s simplicity may support its use as an analytical and communications tool under real-world conditions. However, effective use requires considerable expertise with space technology systems to construct relevant element costs, identify opportunities for efficient shared infrastructure, and assess whether candidate system architectures will meet public objectives. This expertise is generally not economists’ comparative advantage, but is often present within public space agencies. Productive use of this framework therefore requires an iterative process between economists, engineers, and planners at public space agencies, with extensive engagement with industry to obtain parameter estimates and institutional detail.
@@ 4 Future Research Directions
This framework may be most useful as a way to formalize, consolidate, and communicate the various technical and economic assumptions that planners in public space agencies and in industry have adopted. The exercises I have conducted here hopefully help economists see the kinds of insights public space agency planners may be able to use, while also helping public space agency planners see fundamental economic issues more clearly. Making such frameworks practical requires better capabilities in several areas.
4.1 Measuring the Space Economy
Accurate measurement of space economy size and growth is essential for rational planning and investment decisions. Industry associations and investment banks project dramatic growth, with some forecasting markets reaching \$1.8 trillion by 2035 (Daswani et al., 2022; Khlystov et al., 2024). Independent analyses reveal these kinds of estimates often systematically overstate growth, e.g., by double-counting government revenues and input costs (Crane et al., 2020). Despite these measurement problems, academic research often relies on industry projections that assume input costs have fallen dramatically and will continue declining and output markets are competitive, implicitly or explicitly assuming competition in space-related goods and services will continue increasing (Rao et al., 2020; Adilov et al., 2022; Nozawa et al., 2023; Stuermer et al., 2023; Metzger, 2023; Bongers et al., 2025).
These assumptions contradict both economic theory and emerging empirical evidence about space markets. High fixed costs, limited market demand, non-market motivations, and substitutability with terrestrial alternatives constrain entry and long-run competition (Triezenberg et al., 2020; Guyot et al., 2023). Supply constraints persist despite market development efforts (Triezenberg et al., 2024; Rao and Colvin, 2025), launch prices for some government customers have increased in real terms (Kim, 2025), and U.S. space manufacturing capacity utilization has remained flat at about 63% since 2016 (Highfill and Rao, 2025). While real growth in U.S. space manufacturing has exceeded overall U.S. space economy growth (Highfill and Weinzierl, 2024), the persistence of supply constraints suggests structural economic factors may matter more than space-specific policy choices. Better measurement of the space economy can help planners develop better assessments of current conditions, evaluate past market development efforts, and assess the plausibility of projected growth scenarios.
4.2 Estimating Demand for Advanced Technology Systems
In sectors like space launch, statistical estimates of demand can be constructed from historical data (Triezenberg et al., 2020; MITRE, 2021; Adilov et al., 2022). This is less feasible for markets like lunar sample return or lunar surface power, where necessary capabilities do not currently exist. Yet as the framework developed here shows, market demand estimates are critical to designing successful economic architectures.
Developing principled estimates of the scale and elasticity of demand over longer timescales (e.g., 15+ years) for markets that don’t yet exist and have few close analogs—like lunar surface power—is challenging. A structural economic model may be able to measure this demand, e.g., by varying engineering parameters to estimate the total benefit derived from the good or service as in Colvin et al. (2020b) and Metzger (2023), though the relevant preference and technology primitives often require technical knowledge that economists don’t possess. While advances in theory may enable better modeling of such markets, survey and qualitative research methods targeted at relevant stakeholders may provide sufficient institutional detail to usefully constrain models and parameter values. Developing transparent and principled demand and willingness-to-pay estimates for such systems, and communicating them effectively to public space agency planners, may significantly improve the quality of economic architectures for space technology systems.
4.3 Estimating Cost Differentials Across Economic Architectures
Effective use of this framework requires accurate assessment of cost differentials across economic architectures. I assumed both economic architectures had identical costs; this is unlikely in practice. In particular, there is a widespread perception in the space sector that firm fixed-price (FFP) contracts or private development result in lower costs than cost-plus fixed fee (CPFF) contracts or government oversight. These beliefs underlie advocacy for FFP contracting (Vozoff, 2009; Murray, 2023; Berger, 2023) and assumptions that industry-led projects will cost less than government-led projects. [^14] For example, Crane et al. (2017) assumed industry-led projects cost 30% less than equivalent government-led projects in estimating the profitability of private space stations. This assumption was conservative in the context of the analysis it served, but not a neutral one for economic architecture assessments: it would tend to make direct purchases and transfers systematically appear more cost-effective than shared infrastructure investments.
@@ 4.3 Estimating Cost Differentials Across Economic Architectures
Empirical evidence challenges these beliefs. Kim (2025) found NASA has paid increasing real costs for launch services despite using FFP contracts with providers competing in markets where NASA is one among many customers. Carril and Duggan (2020) found the U.S. Department of Defense’s transition from FFP to CPFF contracts during a period of industry consolidation did not statistically significantly increase costs. Kim and Hyde (2025) found industry-led projects show statistically significant cost advantages over government-led projects only for low-risk systems like small satellites. Indeed, simple logic suggests FFP contracts may create “loser’s curse” effects: if contractors can exit when costs exceed expectations, FFP mechanisms transfer upside to contractors while leaving agencies exposed to downside risk. [^15] This is not just a theoretical concern. In 2022, NASA awarded Collins Aerospace a FFP contract under the xEVAS program to develop and produce space suits for ISS missions. The contract would have involved NASA leasing rather than purchasing the suits from Collins, with Collins retaining ownership and the ability to offer the suits to other customers, such as crew on future private space stations (Foust, 2022). In 2024, Collins withdrew from the contract (Foust, 2024). Industry sources suggested that the company faced a combination of cost and schedule overruns and had determined that continued work on the contract was unprofitable. It is unclear what kind of market demand forecasts were used by NASA and Collins when developing the economic architecture for the xEVAS program, and how Collins’ view of the space suits market evolved afterwards. Some observers have speculated that Boeing’s Starliner project may face a similar fate if Boeing decides the LEO crew transportation market is not worth the marginal development effort required (Terlep and Maidenberg, 2025). Given how thoroughly the space community has internalized beliefs about contract type and governance effects on costs, credible econometric analysis could significantly improve acquisition strategies.
4.4 The Relative Value of Government and Market Demand to Private Investors
@@
I have assumed that all revenues, whether derived from government direct transfers, government direct purchases, or market demand, are viewed identically by investors in space technology systems. Is this a reasonable assumption? How do investor perceptions of space-derived cashflows vary with market and government budget volatility? Analyses like event studies of public space company stock prices around contract announcements, surveys of private space investors about their investment evaluation criteria, and choice experiments to elicit investor preferences about uncertain revenue streams could provide quantitative estimates of any systematic differences in how investors in space technology system perceive different revenue sources.
4.5 Analyzing Economic Architectures for Other Space Applications
@@
I have illustrated how principles of economic architecture can help shape programs to develop markets for crewed space services. Triezenberg et al. (2020, 2024) provide detailed analyses of similar goals and trade-offs in the U.S. National Security Space launch market, showing how acquisition decisions affect the number of sustainable competitors. Economic architecture principles can also be applied to orbital debris risk reduction—one of the most studied aspects of space economics in recent years. While grab-and-remove active debris removal (ADR) systems can be cost-ineffective at reducing risk (Colvin et al., 2023; Locke et al., 2024), governments may want to support grab-and-remove technologies for non-economic reasons—such capabilities involve intermediate technologies with military applications, like grappling and relocating non-cooperative objects.
Ultimately, space sustainability programs may require combinations of technologies best provided by different types of actors. Rao et al. (2025) assessed the economic efficiency of portfolios of space sustainability investments including actions and technologies to mitigate risk (e.g., additional shielding against small debris), to remediate risk (e.g., grabbing and removing large debris), and to improve tracking capabilities (e.g., additional radar systems). They identified a class of efficient portfolios containing grab-and-remove ADR that also contain ground-based lasers to remove small debris, new ground-based tracking systems, and improved shielding. Public ownership may be necessary for ground-based laser systems that remove small debris and nudge large objects away from collision trajectories, both due to non-economic factors and because these systems provide non-rival benefits to all satellite operators. Similar considerations may also apply to some ground-based tracking systems that provide baseline situational awareness of the space environment. Direct investments and purchases could seed and support private firms developing and offering ADR services or shielding to satellite operators. The economic architecture questions mirror those in the space station case: determining budget allocation, assessing market demand for debris removal services and shielding, evaluating the distribution of collision risks across satellite operators, and accounting for non-economic factors that influence technology deployment decisions.

### v2

@@ 3 An application to crewed space stations
To estimate the costs of these two architectures, cost data is derived from comparable space station components currently under development or currently operational. Cost estimates for core module elements come from NASA’s Gateway program, which provides the closest available analogs for long-duration crewed space infrastructure: the Power and Propulsion Element (PPE) contract provides cost data for power and thermal systems, the Habitation and Logistics Outpost (HALO) contract provides data for pressurized volume and life support integration, and the Canadian Space Agency’s (CSA) Canadarm3 contract provides data for robotic capabilities. Habitat module construction costs derive from the European Space Agency’s (ESA) Columbus laboratory module, while operating costs are estimated using Columbus’ share of total ISS habitable volume applied to current ISS operating expenses. All cost figures are adjusted to 2025 dollars using NASA’s New Start Inflation Index, which accounts for space technology-specific cost escalation. [^12] Several cost categories are excluded from this analysis. Insurance costs, which can be significant for crewed systems, are omitted due to indications that private insurance markets lack capacity to cover fully privately owned crewed space stations (MacDonald et al., 2024). Launch costs for initial system deployment are excluded, though including them would favor the shared core architecture. Training costs and supply-side structure details for crew operations are abstracted away, though transportation monopolies or inadequate training facilities could complicate the economics of human spaceflight. Integration costs for the core module are assumed to be 15% of total element costs. Figure 2 shows an example of a core module, Figure 3 shows an example of a robotic arm, and Figure 4 shows an example of a habitat module.
@@
Table 2 shows a list of core elements comprising a core module (i.e., foundational infrastructure) as described and, by analogy to roughly similar systems, approximate costs. Table 3 shows the costs for a habitat module (i.e., enabling infrastructure). All figures are expressed in millions of 2025 U.S. dollars and rounded to the nearest million.
Table 2: Core module element cost breakdown. All values rounded to the nearest million.
@@
Table 3: Habitat module cost breakdown. All values rounded to the nearest million.
@@
Operations cost \$232 M/year Columbus has total volume of 75 m3 (ESA, 2025b). ISS has total habitable volume of 388 m3 (NASA, 2025b). 2021 NASA OIG found ISS operating expenses steady at roughly \$1.2B/year over 2016-2020 (NASA, 2021b). FY2026 NASA President’s Budget Request projects similar costs from 2025-2030 (NASA, 2025a). Estimated as ISS total operating expenses scaled by Columbus’ share of habitable volume: $(75/388)\times 1200$.
Habitat module total cost annuity \$418 M/year
These cost estimates allow direct application of the Section 2 framework to assess how the choice between independent stations and shared infrastructure affects competitive viability. Both architectures face identical expected economic conditions: \$500M annually in external demand split between competitors, and \$1B in total capturable program budget. The architectures differ in how this budget gets allocated between direct purchases from firms versus investment in shared infrastructure. In the independent free-flyer architecture, the program budget flows entirely to direct purchases (\$500M per firm assuming two competitors). In the shared core architecture, the program invests in infrastructure—life support, power, docking, and robotic capabilities—that costs \$509M annually based on the component estimates in Table 2, leaving \$491M for direct purchases of habitat services (\$246M per firm). This reflects the functional equivalence of core capabilities across architectures. The shared core provides the same life support, power, and docking functions that firms would otherwise develop independently. This shared requirement is what makes core functions foundational and therefore a candidate for non-rival public provision. Firms are assumed to value the shared core functions identically whether provided by public infrastructure or developed independently. Under this assumption, firms in the shared core architecture benefit from \$509M in cost reduction while receiving \$246M in direct agency purchases. [^13] The \$509M annual value represents what firms would need to spend to develop equivalent capabilities, not necessarily the agency’s actual financing costs, which might differ due to existing assets, different financing terms, or economies of scale in procurement. Table 4 presents the resulting economic parameters for both architectures, while Figure 5 maps these onto the diagram to show how well each approach can sustain competition.
Table 4: Economic parameters under both architectures. All values rounded to the nearest million.
@@
External revenue ($R^{M}_{i}$) \$250 M/year/firm \$250 M/year/firm
@@
Table 4 shows that under the assumed conditions, the independent free-flyer architecture generates negative industry profits of \$355M annually; two competing firms cannot both earn positive returns. Individual firms in this scenario face \$927M in annual costs while receiving only \$750M in combined program and external revenues (\$500M from program-driven purchases plus \$250M from external sources), resulting in \$177M annual losses per firm. The shared core architecture, by contrast, generates positive industry profits of \$154M annually, with each firm earning \$77M in annual profits after accounting for the \$509M cost reduction from sharing foundational infrastructure. Figure 5 maps these results onto the sustainable competition diagram. The free-flyer architecture falls into the region where only monopoly provision is economically viable, while the shared core architecture can sustain two firms. That is, under the assumed conditions, a crewed free-flyer architecture is not compatible with competitive supply of habitat functions. Under the shared core approach firms earn positive profits, though not enough to incent entry of additional competitors. The choice of architecture thus determines whether competition to provide habitat functions is even possible, and the extent to which it may emerge.
@@ 3.1 Exploring alternative assumptions
The baseline analysis demonstrates how shared infrastructure can enable competition under a given set of conditions. The diagram can also be used to explore how different assumptions affect a program’s economics. Two economic variables deserve particular attention because they lie largely outside space agency planners’ control yet can significantly influence market development prospects: the scale of external demand for space station services, and the return expectations of private investors. External demand includes the broader commercial ecosystem for space-based research, manufacturing, and other applications—markets that may grow rapidly if early demonstrations prove successful, or remain limited if technical or economic barriers persist. Investor return expectations, meanwhile, fluctuate with broader macroeconomic conditions including interest rates, risk appetites, and alternative investment opportunities. By examining how the sustainable competition diagram changes under alternative values for these parameters, planners can assess the robustness of different economic architectures and identify conditions under which architectures might succeed.
@@
Figure 6 shows how doubling external demand from \$500M to \$1B annually affects both architectures. This scenario reflects an optimistic case where early space station demonstrations catalyze broader commercial adoption across research institutions, pharmaceutical companies, manufacturing firms, and other potential users. Under these conditions, both architectures become more viable and the choice of architectures becomes less significant. The independent free-flyer architecture, which generated \$355M in annual losses under baseline conditions, now produces positive industry profits and can sustain two competitors with margin for reinvestment or dividends. Each firm in the free-flyer scenario would earn \$500M annually from non-government customers plus \$500M from government purchases, covering the \$927M total costs with \$73M profit margin per firm. The shared core architecture performs even better, generating sufficient profits to sustain three competing firms rather than the two supported under baseline demand, with enough profits to provide returns to existing investors and potentially incent entry of additional competitors. However, new entrants would not be able to enter on the same terms as incumbents without more capacity on the shared core or a new core. This raises an important consideration in using shared infrastructure approaches: the design of the infrastructure may also impose limits on the number of firms that can be sustained under favorable economic conditions. Further analysis of this scenario can identify threshold levels of external demand such that direct purchases alone may sustain competition without requiring shared infrastructure. This outcome depends on successfully developing the commercial ecosystem—an uncertain prospect that may take years or decades to materialize, if ever.
@@
Figure 6: Sustainable competition diagram showing the relationship between industry-level direct government purchases, industry-level total costs, and the potential number of profitable firms for the crewed space stations example with external demand of \$1000 M/year. Shaded regions indicate the expected sustainable number of competitors under different system cost and government-derived revenues.
@@
Figure 7 shows how doubling investor return expectations from 5% to 10% annually affects both architectures. This scenario reflects tighter macroeconomic conditions where rising interest rates, increased risk premiums, or competition from alternative investments raise the hurdle rate for space technology projects. Under 10% return requirements, neither architecture can sustain two competing firms. The independent free-flyer architecture, already unprofitable for two firms under baseline conditions, becomes even less viable with higher costs of capital. The shared core architecture, which supported two firms under baseline conditions, also falls into the monopoly-only region. If there were two firms in the shared core architecture, both would require higher annual profits to compensate investors for the increased return requirement. But revenues and infrastructure-driven cost reductions remain unchanged, and expected cashflows that allowed firms to raise capital under baseline conditions no longer appear as attractive. While space agencies are limited in their abilities to influence macroeconomic conditions, they can monitor them to proactively assess external headwinds or tailwinds and design programs accordingly. Agency-led efforts to aggregate and secure demand from other government entities may be particularly valuable when investors have many other lucrative-seeming opportunities to deploy large amounts of capital and non-government demand for services from habitat functions is weak.
@@
These scenarios reveal several practical insights for public space agencies seeking to develop space markets. First, the choice between shared foundational infrastructure and direct purchases can depend on economic conditions that are outside the agency’s control. Under favorable conditions—e.g., strong external demand and low capital costs—direct purchases may suffice to sustain competition without requiring shared infrastructure investments. Under challenging conditions—e.g., limited external demand or tight capital markets—even well-designed shared infrastructure may prove insufficient to maintain competitive supply. Partnerships with other space agencies to aggregate and channel demand are a potential mechanism to offset challenging external conditions. Second, given how external economic factors can dominate program design choices in determining market outcomes, timing market development efforts to coincide with favorable economic conditions may be as important as program design choices. Space agencies may benefit from integrating macroeconomic assessments into the design of programs that rely on private investors. Third, the diagram provides a systematic way to assess these trade-offs and communicate them to stakeholders, but depends on credible estimates of costs, external demand, and competitive dynamics that remain challenging to develop for novel space technology systems. The estimates may evolve with technical and economic conditions, requiring ongoing assessment and planning.
4 Limitations and Directions for Future Research
The diagram developed here provides a systematic approach to assessing economic trade-offs and uncertainty facing planners designing programs with space market development objectives. Effective application requires expertise with space technology systems to construct relevant element costs, identify opportunities for efficient shared infrastructure, and assess whether candidate system architectures will meet public objectives. This analysis assumes firms are symmetric in costs and capabilities; allowing for asymmetry introduces additional modeling choices such as the order of entry or exit, quickly exploding the number of possible cases. Institutional knowledge can guide analysts toward particular asymmetric scenarios of interest.
@@ 4 Limitations and Directions for Future Research
The diagram may prove most useful as a way to formalize, consolidate, and communicate the various technical and economic assumptions that planners in public space agencies and in industry have adopted. The exercises conducted here may help economists see the kinds of insights public space agency planners may be able to use, while also helping public space agency planners see fundamental economic issues more clearly. Several research directions would enhance the practical applicability of economic analysis for advanced technology markets.
@@
Cost differentials across contract types and governance structures. Effective program design requires accurate assessment of cost differentials across economic architectures. This analysis assumed both economic architectures under consideration had identical costs. However, some in the space sector argue that firm fixed-price (FFP) contracts or private development result in lower costs than cost-plus fixed fee (CPFF) contracts (Berger, 2023). Empirical evidence complicates these arguments. Kim (2025) found NASA has paid increasing real costs for launch services despite using FFP contracts with providers competing in markets where NASA is one among many customers. Carril and Duggan (2020) found the U.S. Department of Defense’s transition from FFP to CPFF contracts during a period of industry consolidation did not statistically significantly increase costs. Kim and Hyde (2025) found industry-led projects show statistically significant cost advantages over government-led projects only for low-risk systems like small satellites. FFP contracts may also create “loser’s curse” effects: if contractors can exit when costs exceed expectations, FFP mechanisms transfer upside to contractors while leaving agencies exposed to downside risk. [^14] For example, in 2022 NASA awarded Collins Aerospace a FFP contract under the xEVAS program to develop and produce space suits for ISS missions. Collins withdrew from the contract in 2024, reportedly due to cost and schedule overruns (Foust, 2022, 2024). Some observers have speculated that Boeing’s Starliner project may face a similar fate (Terlep and Maidenberg, 2025). Credible econometric analysis of contract type and governance effects on costs could improve acquisition strategies and economic architecture assessments.
@@
Investor perceptions of different revenue sources. This analysis assumed that investors view all revenues, whether derived from a government or non-government source, identically. Is this a reasonable assumption? How do investor perceptions of space-derived cashflows actually vary with market and government budget volatility? Do investors systematically value government contracts differently from commercial revenues? Analyses like event studies of public space company stock prices around contract announcements, surveys of private space investors about their investment evaluation criteria, and choice experiments to elicit investor preferences about uncertain revenue streams could provide quantitative estimates of any systematic differences in how investors perceive different revenue sources. Such evidence would allow more accurate modeling of capital costs and return requirements across different economic architectures.
External demand forecasting. Assessing economic trade-offs in program design requires estimates of external demand. For nascent markets like commercial space stations, external demand is deeply uncertain across multiple dimensions, e.g., how much demand will materialize, what services customers will demand, and when they will demand them. In the near term, external demand for advanced space systems may consist primarily of other government entities whose budgets and priorities the focal agency cannot control. Over longer timescales, the scale and composition of private commercial demand remains uncertain. Different customer types demand different service profiles: government agencies typically prioritize long-duration missions that signal capability and prestige, while commercial customers may prefer shorter-duration access for specific applications. Improved methods for forecasting demand under deep uncertainty—including both the scale of demand from various sources and the nature of services demanded—would help planners design better market development programs.
@@
Applications to other domains. This kind of economic analysis can be applied beyond crewed space stations. Triezenberg et al. (2020, 2024) provide detailed analyses of similar goals and trade-offs in the U.S. National Security Space launch market, showing how acquisition decisions affect the number of sustainable competitors. Orbital debris risk reduction programs offer another example. While grab-and-remove active debris removal (ADR) systems can be cost-ineffective at reducing risk (Colvin et al., 2023; Locke et al., 2024), governments may want to support such technologies for non-economic reasons—these capabilities involve intermediate technologies with potential military applications. Rao et al. (2025) assessed the economic efficiency of portfolios of space sustainability investments, identifying efficient combinations that include grab-and-remove ADR alongside ground-based lasers for small debris removal, new tracking systems, and improved shielding. Public ownership may be necessary for ground-based laser systems and some tracking systems that provide non-rival benefits to all satellite operators. Direct investments and purchases could support private firms developing ADR services or shielding. The economic architecture issues mirror those in the space station case: budget allocation, external demand assessment, risk distribution, and accounting for relevant non-economic factors.